How am I going to find it's inverse given:3x+5x=10
Found 2 solutions by Edwin McCravy, ikleyn:
Answer by Edwin McCravy(20055)

Did you mean 3x+5x=10 which is just 8x=10 or 4x=5 or x=5/4 which is a vertical
line and not a function?  If so, its inverse is simply found by replacing the
x by a y and getting 4y=5 or y=5/4 which is a horizontal line and which is a
function.

But I'll bet your teacher is not trying to see whether you can add 3x and 5x to
get 8x.  I'll bet instead that you accidentally made a typo and one of those x's
was supposed to be a y.  Am I right?  Check and see.

Edwin

Answer by ikleyn(52781)

.
How am I going to find it's inverse given:3x+5x=10
~~~~~~~~~~~~~~~~~~~~~~~


This "problem" recalls me two-legged horse of Baron Munchausen.

It needs a major repair.
All the other legs need to be attached.
Otherwise it is only fit to be thrown in the trash.